# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ V-D Multi-Link Connectivity
As mentioned earlier, some emerging 6G applications, such as aerial vehicle control and autonomous driving, impose stringent requirements in terms of both latency and connection reliability, a paradigm referred to as URLLC. In this scenario, multi-link connectivity (or MC) has shown the potential to meet the URLLC service requirements [53]. Specifically, MC architectures can be categorized at the BS level (i.e., the user connects to multiple BSs within the same network, usually managed by the same operator), at the network level (i.e., the user is connected to different networks, possibly from different operators), or at the technology level (i.e., the user connects via different RATs) [54]. In this paper, we mostly focus on the first category, which also includes techniques such as coordinated multi-point (CoMP), which gained significant importance in LTE-advanced and 5G, where multiple BSs within the same network simultaneously serve the same user.
V-D1 MC at the BS level
In [145], the authors illustrated the benefits of MC for latency-sensitive applications like video streaming over local wireless network. However, the system model is limited to a local network scale, and cannot support high-speed transportation systems, such as high-speed trains. To fill this gap, the study [146] focused on dynamic multi-link scheduling for cellular networks serving high-speed train passengers. The study aimed at optimizing the use of existing public cellular network resources to provide more reliable and stable connections by minimizing latency and packet loss rates. Their findings demonstrated improvements in transmission time and bandwidth compared to various baseline methods.
In [147], air-to-air communication was studied using a 3D model integrating aerial BSs, aerial UEs, TNs, and GUEs. Two UAV types were considered: aerial BSs connecting via HAPS/TBS, and aerial users relaying data to GUEs. A joint-transmission CoMP scheme was proposed where each aerial user is served by four aerial BSs forming a tetrahedral cell. The study derived data rates and coverage probabilities, and introduced a frequency allocation scheme to reduce inter-cell interference. Compared to the binomial-Voronoi dynamic cooperation set, the approach achieved similar performance with much lower complexity and signaling overhead.
In [148], aerial UE connectivity was studied using CoMP transmissions for UAVs served by TBSs. Two scenarios were analyzed—static hovering and mobile UAVs—with coverage probability bounds derived and handover rates evaluated via a custom mobility model. Results showed that CoMP transmission greatly improves achievable coverage probability.
In [149], an uplink CoMP scenario was considered where UAVs collect and forward signals from GUEs. A UAV placement and movement scheme was proposed to maximize average throughput for mobile GUEs, though UAV energy constraints remain a critical challenge.
In [54], an MC SAGIN algorithm was proposed for remote aerial vehicle (AV) piloting beyond LoS. The system model included LEOs, HAPS, and TBSs for AV communication, along with direct inter-AV links. However, it excluded LEO-HAPS and LEO/HAPS-TBS links, assuming both LEO and HAPS connect to the core network via their ground stations.
The transmission strategy in [54] used message cloning, sending multiple copies via all available links for independent decoding at the AV. The model incorporated practical antenna setups, unreliable backhaul, buffer queues, and fading effects. Reliability and availability of different MC configurations were evaluated through simulations, with brute-force optimization used to find the optimal MC path ensuring QoS. Results showed that LoS interference strongly impacts link performance, and MC is vital for safe operation of certain AVs.
A latency-sensitive scenario was explored in [150], which addressed the issue of MC computational task offloading in NTN. The system model featured a HAPS and several UAVs providing on-demand MEC services to randomly distributed GUEs. The simulation results showed that MC offloading strategies can significantly reduce task computation delay.
V-D2 MC at the network level
In [151], network-level MC was studied for smart agriculture, where users simultaneously connected to satellite and terrestrial networks. Latency was the key metric, evaluated via drive tests using Starlink and two public 5G non-standalone TBSs. Results showed that MC greatly reduced latency compared to single connectivity with either satellite or TNs.
In [152], latency in an MC scenario with a UAV-based client was measured across four flight heights, comparing a single-operator setup with a dual-operator hybrid access system, which improved latency and reliability toward 3GPP targets. Similarly, [153] evaluated MC across two public LTE networks in urban, suburban, and rural settings with a UAV-based UE, showing that operator-specific performance variations make network-level MC effective for enhancing reliability.
V-E Mobility Management and Handover
Mobility management faces major challenges in seamless handovers, especially for LEO and HAPS systems. In [154], the standard 5G NR handover procedure was evaluated for a high-speed train in a LEO-based deployment versus a 3GPP urban macro TN. Simulations showed radio connection failures during handovers were 10× higher and outages 5× longer in the NTN scenario.
The authors in [155] proposed a platform that emulates hybrid satellite–terrestrial networks, to simulate TN/NTN handover procedures. This approach overcomes the need to physically launching satellites, and provides a cost-effective time-efficient simulation environment prior to real-world implementation.
In [156], three inter-satellite handover methods were compared for space-to-ground performance: Closest Satellite, Max Visibility, and SINR-based. The first two improved spectral efficiency and reduced delays, while the SINR-based approach minimized the number of handovers.
In [157], handover performance in a 5G NR network over LEO satellites with Earth-moving cells was studied. An intra-satellite mobility solution was proposed that avoids UE radio measurements by exploiting predictable satellite trajectories and antenna gains. Simulations in rural and urban scenarios showed reduced service failures and unnecessary handovers, with improved downlink SINR.
In [158], Event A4 was proposed for mobility management in LEO satellite networks, since minimal RSRP variation in NTNs makes A3-based handovers ineffective. Unlike A3, which compares serving and neighboring cells, A4 triggers when RSRP exceeds an absolute threshold. Simulations showed A4-based handovers outperform A3 in reliability and robustness, though determining cell borders remains difficult in NTNs, especially for fast-moving UEs.
In [159], an angle-based handover method was proposed for LEO networks, where the satellite elevation angle is calculated from its separation with the UEs and cell center. This approach was shown to provide reliable, continuous connectivity for UAVs in remote and challenging areas.
In [160], a graph theory-based algorithm was proposed for UAV handovers between terrestrial and satellite networks. By adapting in real time to UAV mobility and network changes, it minimizes latency and disruptions. The algorithm considers RSRP, signal to noise ratio (SNR), and elevation angle to optimize handover timing and ensure smooth connectivity.
The authors in [37] developed various handover techniques for an integrated LEO-HAPS network, and evaluated the handover delay and signaling overhead. Specifically, they formulated the problem of determining the optimal handover triggering time, which was analytically solved using Lagrange duality.
In [161], a mobility management solution was proposed for integrated HAPS-ground networks, allowing aerial users to connect to either terrestrial BSs or a HAPS. The approach focused on minimizing unnecessary handovers while ensuring service continuity and quality of experience.
In [162], an adaptive handover scheme with cooperative transmissions was proposed for HAPS systems to mitigate outage probability and long service interruptions. The algorithm adaptively selects the platform with the highest channel gain based on user mobility and real-time channel conditions, improving reliability and service continuity.
In [163], the impact of HAPS vertical and swing movements on handover performance was analyzed. A refined ground coverage model and analytical expressions were developed, showing that path loss strongly affects coverage, with swing movement exerting a greater influence on handover probability than vertical movement.
In [164], an adaptive handover algorithm was proposed for quasi-stationary HAPS, addressing the limitations of fixed-threshold methods. Using a least mean square prediction-based power model, the scheme adjusts handover parameters according to terminal speed and the confidence levels of expected signal strength, improving reliability under HAPS mobility conditions.
Furthermore, a new handover approach based on convolutional neural networks was introduced in [165]. The authors modeled the handover process as a directed graph, where a user with (partial) knowledge of its future signal strength can identify the optimal handover sequence. The CNN was trained to learn near-optimal handover strategies from historical signal strength data.
V-F Antenna Technologies
Integrating advanced antenna technologies, such as massive multiple-input multiple-output (mMIMO), spatial multiplexing, digital beamforming, and RISs, into HAPS systems can improve network capacity and efficiency across a range of diverse and challenging environments.
V-F1 MIMO-like antennas
The authors in [166] proposed a cylindrical mMIMO system for HAPS. They showed that the cylindrical mMIMO shape provides up to 2.1 times higher capacity compared to traditional planar arrays, with an extended coverage of up to 100 km in radius, and improves the SINR by approximately 10 dB at the cell edge.
In [167], a vertical uniform linear array sector-based cylindrical antenna was proposed for HAPS, combined with a NOMA clustering method to mitigate strong geographic user correlation. An optimization algorithm for power distribution was introduced to maximize energy efficiency, spectral efficiency, and QoS using successive interference cancellation. Simulations showed that spatial correlation strongly affects spectrum and energy efficiency in multi-antenna HAPS systems.
In [168], a two-stage precoding system for TN-HAPS spectrum sharing was proposed using phased array antennas. The first stage forms nulls toward TN cells, and the second directs beams to HAPS users. With a 196-element cylindrical array, the approach improved SINR for TN users by up to 10 dB over a conventional minimum mean square error (MMSE) precoder.
In [169], RF precoder design was formulated as an average signal to leakage interference and noise ratio (SLNR) maximization problem using a reduced-column discrete Fourier transform (DFT) matrix. The derived lower bound was converted into a trace problem for initial precoder design, and a greedy algorithm was proposed to update the RF precoder, achieving superior performance over existing methods.
@@ V-F1 MIMO-like antennas
In [170], a hemispherical antenna array for HAPS was proposed to maximize gain by aligning users with specific antenna elements. Using analog beamforming, user data rates were estimated from steering vectors, while an antenna selection algorithm and Bisection-based power allocation optimized performance. Results showed cumulative data rates up to 14 Gbps.
The authors in [99] proposed an algorithm to mitigate inter-cell interference between GUEs and neighboring UAV-BSs. The scheme operates in two stages: GUEs send messages to UAVs in the sub-6 GHz band, and UAVs then allocate resources and beamform signals to a HAPS in the sub-THz band acting as a CPU. An optimization problem was formulated to maximize the minimum SINR by optimizing power allocation and UAV positions.
In [171], multi-beam antenna technology was used to enhance coverage by adaptively adjusting beam widths within elliptical cells, equalizing signal quality and reducing the number of cells while maintaining coverage and link quality. The study also showed that integrating multi-beamforming with space-time block coding can further increase capacity without extra bandwidth.
The authors in [172] evaluated the advantages of using a triangular (rather than square) lattice for phased arrays in HAPS systems operating in the 5G NR n257 and n258 frequency bands. By considering a 64-element planar array, the results demonstrated that the triangular lattice array offers superior performance in terms of array gain, average sidelobe level, and mutual coupling, making it a more effective choice for 5G mMIMO systems.
V-F2 RIS antennas
The use of RISs can enhance signal quality, coverage, and LoS connectivity at low cost, and when combined with mMIMO, improve data rates and reliability [173]. In [174], a HAPS-RIS system was studied where RISs serve as amplify-and-forward relays to bypass blockages. Evaluation via BER, capacity, and outage probability showed significant performance gains, confirming the potential of integrating RISs with HAPS.
Another example of a HAPS-RIS system was considered in [175]. The RIS was mounted on the HAPS side for reflecting signals from GUEs to an upstream satellite. The authors used DRL for the joint optimization of RIS phase shifts and transmit beamforming, which significantly improved system efficiency.
In [176], a RIS was installed on a HAPS to provide extensive LoS and full-area coverage for UAV-BSs deployed in response to sudden urban traffic spikes. The combined optimization of RIS phases, array-partitioning algorithms, and positioning enhanced the energy efficiency of each UAV-BS.
In [177], an energy-optimization algorithm was proposed for HAPS, enabling mode switching to reduce consumption and extend loitering duration. The modes include a low-power HAPS-RIS mode for passive communication and a HAPS-SMBS mode for enhanced computing, caching, and communication. The system prioritizes the passive mode, activating higher-energy modes only when additional capacity is required.
V-G Channel Estimation
As 6G moves to high-frequency bands (cmWave, mmWave, sub-THz), channel conditions worsen due to path loss, fading, and interference, making accurate estimation essential for reliable connectivity and low latency. Advanced techniques such as beamforming, mMIMO, and ISAC also depend on precise CSI. In [178], channel estimation errors in an integrated satellite-HAPS-terrestrial system were analyzed, and a NOMA-based power allocation scheme was proposed to meet QoS requirements.
In [179], a HAPS-LEO architecture was studied to address channel estimation and synchronization challenges. A CNN-based estimator was proposed to mitigate carrier frequency offset from residual Doppler effects, with results showing improved mean square error and BER performance over benchmark methods.
The study in [180] evaluated LTE downlink performance over HAPS, comparing least square and MMSE channel estimation under a Ricean fading model, and found that GUEs at low elevation angles suffer degradation from multipath and Doppler. Similarly, [181] analyzed a single-carrier LTE HAPS channel, showing that elevation angles above 40° yield optimal conditions, with performance influenced by bandwidth, modulation, and Doppler shifts. More recently, [182] compared least square, MMSE, and MMSE+DFT estimation methods for HAPS-based OFDM systems.
In [183], sparse HAPS channels were estimated using a superimposed pilot sequence method that leverages prior knowledge of UEs’ locations. The model featured a HAPS-based beamforming multi-antenna BS serving mobile GUEs, with backhaul via ground stations and satellites. Numerical results examined the influence of antenna configuration, channel sparsity, SNR, and interferer density.
In [184], a VHetNet combining GEO/LEO satellites, HAPS, and TN was studied with full-duplex RIS-assisted backhauling over HAPS channels. A graph attention network-based channel estimator was proposed to address channel variability and hardware impairments, requiring estimation at only one transmitting node without on-off control or half-duplex switching. Simulations confirmed strong performance across diverse channel and hardware conditions.
V-H Interference Management and Beamforming
In [185], stratosphere-to-Earth co-channel interference from HAPS was analyzed, with a methodology to assess its impact on TNs through performance degradation. The study accounted for HAPS mobility, realistic azimuth/elevation distributions of terrestrial microwave links, and progressive HAPS loading, providing a comprehensive framework for interference evaluation.
The study [186] examined a 3D mMIMO system where a HAPS mounting a horizontal planar antenna array hovers in the sky, serving multiple GUEs. The study modeled intra-cell uplink interference using Beta distributions, and demonstrated that adjusting the altitude of the array can significantly reduce interference, thereby improving key performance metrics such as signal to interference ratio, coverage, and throughput.
In [94], interference between terrestrial and satellite platforms was studied, and a beamforming method using ephemeris data was proposed to dynamically suppress interference. By creating spatial nulls toward visible satellites, the technique protected satellite uplinks while maintaining terrestrial downlink performance.
To expand the operational coverage area for UAVs in a segregated airspace, Ku-band links can be utilized. In [187], the authors proposed a new conformal phased array antenna, developed with ultra-lightweight aerogel material, to support this architecture while managing interference. This antenna design offered high performance at low size, weight, and power.

### v2

@@ VI-D Multi-Link Connectivity
In this context, multi-link connectivity (or MC) has shown the potential to meet the URLLC service requirements [187]. Specifically, MC architectures can be categorized at the BS level (i.e., the user connects to multiple BSs within the same network, usually managed by the same operator), at the network level (i.e., the user is connected to different networks, possibly from different operators), or at the technology level (i.e., the user connects via different RATs) [186]. In this paper, we focus on the first category, which also includes techniques such as coordinated multi-point (CoMP), which gained significant importance in LTE-advanced and 5G, where multiple BSs within the same network simultaneously serve the same user.
Li et al. [133] studied air-to-air communication using a 3D model integrating aerial BSs, aerial UEs, TNs, and GUEs. Two UAV types were considered: aerial BSs connecting via HAPS/TBS, and aerial users relaying data to GUEs. A joint-transmission CoMP scheme was proposed where each aerial user is served by four aerial BSs forming a tetrahedral cell. The study derived data rates and coverage probabilities, and introduced a frequency allocation scheme to reduce inter-cell interference. Compared to the binomial-Voronoi dynamic cooperation set, the approach achieved similar performance with much lower complexity and signaling overhead. Meanwhile, Amer, Saad, and Marchetti [25] studied aerial UE connectivity using CoMP transmissions for UAVs served by TBSs. Two scenarios were analyzed—static hovering and mobile UAVs—with coverage probability bounds derived and handover rates evaluated via a custom mobility model. Results showed that CoMP transmission greatly improves achievable coverage probability. Prior to that, Liu, Zhang, and Zhang [136] considered an uplink CoMP scenario where UAVs collect and forward signals from GUEs. A UAV placement and movement scheme was proposed to maximize average throughput for mobile GUEs, though UAV energy constraints remain a critical challenge. Recently, Salehi, Ozger, and Cavdar [186] proposed an MC SAGIN algorithm for remote aerial vehicle (AV) piloting beyond LoS. The system model included LEOs, HAPS, and TBSs for AV communication, along with direct inter-AV links. However, it excluded LEO-HAPS and LEO/HAPS-TBS links, assuming both LEO and HAPS connect to the core network via their ground stations. The transmission strategy in [186] used message cloning, sending multiple copies via all available links for independent decoding at the AV. The model incorporated practical antenna setups, unreliable backhaul, buffer queues, and fading effects. Reliability and availability of different MC configurations were evaluated through simulations, with brute-force optimization used to find the optimal MC path ensuring QoS. Results showed that LoS interference strongly impacts link performance, and MC is vital for safe operation of certain AVs. At thesame time, latency-sensitive scenario was explored by Sadovaya et al. [183], which addressed the issue of MC computational task offloading in NTN. The system model featured a HAPS and several UAVs providing on-demand MEC services to randomly distributed GUEs. The simulation results showed that MC offloading strategies can significantly reduce task computation delay.
VI-E Mobility Management and Handover
Mobility management faces major challenges in seamless handovers, especially for LEO and HAPS systems. The architectural and functional principles of connected-mode mobility and handover (HO) in NR and LTE are specifically described in 3GPP TS 38.300 and TS 36.300 [9, 6], respectively. Referencing the specific handover signaling and timing protocols outlined in the related Radio Resource Control Radio Resource Control (RRC) standards, these specifications specify the functions of radio access network (RAN) nodes, mobility statuses, and high-level handover concepts. In the context of HAPS-assisted mobility, where wide coverage regions and quasi-stationary platforms produce different handover dynamics than terrestrial networks, its applicability to NTN scenarios is examined.
Juan et al. [112] evaluated the standard 5G NR handover procedure for a high-speed train in a LEO-based deployment versus a 3GPP urban macro TN. Simulations showed radio connection failures during handovers were 10× higher and outages 5× longer in the NTN scenario. Further, an intra-satellite mobility solution was proposed [113] that avoids UE radio measurements by exploiting predictable satellite trajectories and antenna gains. Simulations in rural and urban scenarios showed reduced service failures and unnecessary handovers, with improved downlink SINR. Zhang et al. [243] introduced a convolutional neural networks-based handover approach for LEO satellite network. The authors modeled the handover process as a directed graph, where a user with (partial) knowledge of its future signal strength can identify the optimal handover sequence. The CNN was trained to learn near-optimal handover strategies from historical signal strength data. Voicu, Bhattacharya, and Petrova [218] compared three inter-satellite handover methods for space-to-ground performance: Closest Satellite, Max Visibility, and SINR-based. The first two improved spectral efficiency and reduced delays, while the SINR-based approach minimized the number of handovers. Warrier et al. [226] introduced a graph theory-based algorithm for UAV handovers between terrestrial and satellite networks. By adapting in real time to UAV mobility and network changes, it minimizes latency and disruptions. The algorithm considers RSRP, signal to noise ratio (SNR), and elevation angle to optimize handover timing and ensure smooth connectivity.
The quasi-stationary nature of HAPS in the stratosphere introduces additional challenges for conventional handover algorithms, which rely on fixed thresholds. He, Cheng and Ni [91] introduced an adaptive handover algorithm, addressing the limitations of fixed-threshold methods, in order to handle possible HAPS disturbance. Using a least mean square prediction-based power model, the scheme adjusts handover parameters according to terminal speed and the confidence levels of expected signal strength, improving reliability under HAPS mobility conditions. Further [90] a cooperative transmissions scheme was proposed for HAPS systems to mitigate outage probability and long service interruptions. The algorithm adaptively selects the platform with the highest channel gain based on user mobility and real-time channel conditions, improving reliability and service continuity. Later [224] Wang, Li, and Zhou analyzed the impact of HAPS vertical and swing movements on handover performance. A refined ground coverage model and analytical expressions were developed, showing that path loss strongly affects coverage, with swing movement exerting a greater influence on handover probability than vertical movement. Menwhile, Li et al. [131] developed various handover techniques for an integrated LEO-HAPS network, and evaluated the handover delay and signaling overhead. Specifically, they formulated the problem of determining the optimal handover triggering time, which was analytically solved using Lagrange duality. Later, Almuallim et al. [23] studied a mobility management solution for integrated HAPS-ground networks, allowing aerial users to connect to either terrestrial BSs or a HAPS. The approach focused on minimizing unnecessary handovers while ensuring QoS.
VI-F Antenna Technologies
Integrating advanced antenna technologies, such as massive multiple-input multiple-output (mMIMO), spatial multiplexing, digital beamforming, and RISs, into HAPS systems can improve network capacity and efficiency across a range of diverse and challenging environments. For example, Dicandia and Genovesi [60] evaluated the advantages of using a triangular (rather than square) lattice for phased arrays in HAPS systems operating in the 5G NR n257 and n258 frequency bands. By considering a 64-element planar array, the results demonstrated that the triangular lattice array offers superior performance in terms of array gain, average sidelobe level, and mutual coupling, making it a more effective choice for 5G mMIMO systems. Meanwhile, Tashiro, Hoshino, and Nagate [208] proposed a cylindrical mMIMO system for HAPS. They showed that the cylindrical mMIMO shape provides up to 2.1 times higher capacity compared to traditional planar arrays, with an extended coverage of up to 100 km in radius, and improves the SINR by approximately 10 dB at the cell edge. Further, in order to consider the coexistence with terrestrial systems, the authors introduced a two-stage precoding system for TN-HAPS spectrum sharing using phased array antennas [209]. The first stage forms nulls toward TN cells, and the second directs beams to HAPS users. With a 196-element cylindrical array, the approach improved SINR for TN users by up to 10 dB over a conventional minimum mean square error (MMSE) precoder. Cylindrical antenna configuration was later considerd by Shafie et al. [191]. The authors proposed a vertical uniform linear array sector-based cylindrical antenna for HAPS, combined with a NOMA clustering method to mitigate strong geographic user correlation. An optimization algorithm for power distribution was introduced to maximize energy efficiency, spectral efficiency, and QoS using successive interference cancellation. Simulations showed that spatial correlation strongly affects spectrum and energy efficiency in multi-antenna HAPS systems.
@@ VI-F Antenna Technologies
Alternative antenna configuration was studied by Abbasi, Yanikomeroglu, and Kaddoum [12] to address the suboptimal gain for users located beneath the HAPS. The authors introduced a hemispherical antenna array for HAPS to maximize gain by aligning users with specific antenna elements. Using analog beamforming, user data rates were estimated from steering vectors, while an antenna selection algorithm and Bisection-based power allocation optimized performance. Results showed cumulative data rates up to 14 Gbps. The authors also proposed a cell-free scheme to mitigate inter-cell interference between GUEs and neighboring UAV-BSs [14]. The scheme operates in two stages: GUEs send messages to UAVs in the sub-6 GHz band, and UAVs then allocate resources and beamform signals to a HAPS in the sub-THz band acting as a CPU. An optimization problem was formulated to maximize the minimum SINR by optimizing power allocation and UAV positions.
Regardin the UE antenna systems, Kim et al. [122], proposed a dual-polarized shared aperture antenna for Ku- and Ka-bands, developed through a four-step derivation process using slot aperture coupling to achieve broadside radiation in both bands. Based on array factor theory, the optimal inter-element spacing was determined, resolving the trade-off between beam steering angles and array spacing in dual-band systems. The use of RISs is another approach to improve signal quality and coverage. In fact, low-cost RIS components enable customizable communication channels, LoS connectivity, and obstruction avoidance [235]. When combined with traditional mMIMO techniques, RISs can also boost data rates and reliability. In a system where HAPS relay data via RISs due to transmission blockages, the RIS acts as an amplify-and-forward relay between source and ground users. This approach was evaluated by Odeyemi, Owolawi, and Olakanmi [167] through metrics like BER, capacity, and outage probability, with analytical and asymptotic results showing significant performance improvements for HAPS and RISs integration.
VI-G Channel Estimation
6G is expected to use high-frequency bands such as cmWave, mmWave or even sub- THz, to support new scenarios for seamless ubiquitous 3D connectivity (see Sec. V-C). However, the channel conditions in these bands will become more variable, suffering from significant path loss, fading, and interference. Thus, accurate channel estimation is necessary to adapt transmission parameters dynamically, ensure reliable connectivity, and maximize data throughput while minimizing latency. Advanced techniques like beamforming, mMIMO, and ISAC, also require precise channel state information (CSI) via channel estimation, to provide reliable communication services. In this context, Iskandar and Aziz [103] evaluated LTE downlink performance over HAPS, comparing least square and MMSE channel estimation under a Ricean fading model, and found that GUEs at low elevation angles suffer degradation from multipath and Doppler. Later Hidayat and Iskandar [94] analyzed a single-carrier LTE HAPS channel, showing that elevation angles above 40° yield optimal conditions, with performance influenced by bandwidth, modulation, and Doppler shifts. More recently, Rahayu and Iskandar [176] compared least square, MMSE, and MMSE+ discrete Fourier transform (DFT) estimation methods for HAPS-based OFDM systems. Also, Nawaz et al. [164] estimated sparse HAPS channels using a superimposed pilot sequence method that leverages prior knowledge of UEs’ locations. The model featured a HAPS-based beamforming multi-antenna BS serving mobile GUEs, with backhaul via ground stations and satellites. Numerical results examined the influence of antenna configuration, channel sparsity, SNR, and interferer density. Later, Tekbıyık et al. [210] studied a vHetNet combining GEO/LEO satellites, HAPS, and TN with full-duplex RIS-assisted backhauling over HAPS channels. A graph attention network-based channel estimator was proposed to address channel variability and hardware impairments, requiring estimation at only one transmitting node without on-off control or half-duplex switching. Simulations confirmed strong performance across diverse channel and hardware conditions. Meanwhile, Güven and Karabulut Kurt [85] studied a HAPS-LEO architecture to address channel estimation and synchronization challenges. A CNN-based estimator was proposed to mitigate carrier frequency offset from residual Doppler effects, with results showing improved mean square error and BER performance over benchmark methods. And recently, Guo et al. [84] analyzed the effect of channel estimation errors in an integrated satellite-HAPS-terrestrial communication system, and proposed a novel NOMA-based power allocation scheme to satisfy QoS requirements.
VI-H Interference Management and Beamforming
Interference management and beamforming technologies enhance HAPS communication by jointly improving signal strength and reducing interference in remote and demanding areas. A comprehensive analysis of stratosphere-to-Earth co-channel interference caused by HAPS was presented by Milas and Constantinou [156], who introduced a methodology for assessing the impact of this interference component on TNs by evaluating performance degradation. The methodology considered various factors, including the mobility of HAPS, the realistic distribution of azimuth and elevation angles of terrestrial microwave links, and the gradual loading of HAPS structures. A 3D mMIMO system was examined by Hu, Hong and Evans [98], who considered an uplink scenario with pure LoS propagation from multiple GUEs to a HAPS mounting a horizontal planar antenna array. The study modeled intra-cell interference using Beta distributions, and demonstrated that adjusting the altitude of the array can significantly reduce interference, thereby improving key performance metrics such as signal to interference ratio, coverage, and throughput.